OpenStudy (alex_mattucci):

I need help dividing a polynomial!!! (10x4y3 + 5x3y2 - 15x2y - 25x2y4)/-5x2y

Nnesha (nnesha):

Divide the coefficients and for variables apply the exponent rule when we divide same base , `subtract` their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

Nnesha (nnesha):

like for first one \[\huge\rm \frac{ 10 x^4y^3}{ -5x^2y^1} =\frac{10}{-5}x^{4-2}y^{3-1}\] y is same as y to the one power

Nnesha (nnesha):

subtract exponents and divide the coefficient

OpenStudy (alex_mattucci):

Ok,thank you!!!

OpenStudy (alex_mattucci):

Just to be correct on this, That would make the answer -2x^2y^2-1xy+3+5y^2 correct @Nnesha?

Nnesha (nnesha):

\[\huge\rm \frac{ - 25x^2y^4 }{ -5x^2y } = ??\]

Nnesha (nnesha):

remember y is same as y to the one power (y^1)